# GLOBALIZATION? TRADE WAR? A COUNTERBALANCE PERSPECTIVE

### v4

@@ 4.2 The Bargaining Solution
$\frac{1}{q}=\frac{1}{\max\left\{\frac{\pi_{j}}{\pi_{i}},\frac{g_{j}}{g_{i}}\right\}}=\min\left\{\frac{\pi_{i}}{\pi_{j}},\frac{g_{i}}{g_{j}}\right\}.$
We have two scenarios when country $i$ and $j$ have opposite preferences on $p$ and $q$. In the first scenario, $j$ prefers $q$ to $p$, whereas $i$ prefers $1/p$ to $1/q$. So $u_{j}(x)=x$ is an utility function for $j$ because $u_{j}(q)=q>p=u_{j}(p)$. Similarly, $u_{i}(x)=x$ is an utility function for $i$ since $u_{i}(1/p)=1/p>1/q=u_{i}(1/q)$. These are von Neumann-Morgenstern utilities, uniquely determined up to a positive affine transformation. By Eq. (15), $p$ and $1/q$ are their respective status quo utilities, obtained if one decides not to bargain with the other player. Assuming that all countries have equal prior bargaining power, the Nash bargaining solution solves the following maximization problem
$\max\limits_{\lambda_{ji}\lambda_{ij}=1;\ \lambda_{ji}>0}\left(\lambda_{ji}-p\right)\left(\lambda_{ij}-\frac{1}{q}\right).$ (16)
[…]
Figure III: The Nash bargaining solution $\lambda_{ji}^{*}$ for Eq. (16) and (17), respectively.
In the second scenario of opposite preferences, $j$ prefers $p$ to $q$ while $i$ prefers $1/q$ to $1/p$. Their utility functions are $u_{j}(x)=q-x$ and $u_{i}(x)=1/p-x$, respectively, and the status quo is 00 for both. The maximization problem becomes
$\max\limits_{\lambda_{ji}\lambda_{ij}=1;\ \lambda_{ji}>0}\left(q-\lambda_{ji}\right)\left(\frac{1}{p}-\lambda_{ij}\right).$ (17)
The right plot of Figure III illustrates this situation where $j$ moves from $q$ to $p$ and $i$ from $1/p$ to $1/q$ along the curve. They meet at the solution point $(\lambda^{*}_{ji},\lambda_{ij}^{*})$ which maximizes the area of the rectangle, cornered by $(\lambda^{*}_{ji},\lambda^{*}_{ij})$ and $(q,1/p)$. In this utility function, $\lambda_{ji}$ seems to have no meaning of reprisal. However, we can rephrase Eq. (17) by
$\max\limits_{\lambda_{ji}\lambda_{ij}=1;\ \lambda_{ji}>0}\left(\lambda_{ji}-q\right)\left(\lambda_{ij}-\frac{1}{p}\right)$ (18)
in which $\lambda_{ji}-q$ and $\lambda_{ij}-1/p$ are the loss functions for $j$ and $i$, respectively. To coerce an agreement formation, each country plays a threat strategy to maximize its counterpart’s loss. The final outcome is Eq. (18). Moving from $p$ to $q$, a larger $\lambda_{ji}$ means less loss to $j$; thus, $\lambda_{ji}$ still preserves the meaning of retaliation for $j$ in Eq. (18) and, so, in Eq. (17).
Theorem 5 answers the maximization problems, and the solution satisfies the identity $\lambda_{ji}\lambda_{ij}=1$ by design. The solution is the same for both preference scenarios, so it also works when country $i$ or $j$ is indifferent between greater competitiveness and improved trade balance. In Eq. (16), (17) and (18), both countries have a symmetric role and the same prior bargaining power. This is an axiom of the Nash bargaining. The solution remains unchanged under any positive affine transformation of the utility functions, according to another axiom of the bargaining solution. By taking the square root of $\pi_{j}/\pi_{i}$, $\lambda_{ji}^{*}$ mitigates the Matthew effect in $\pi$. If $g_{j}<g_{i}$, for example, then $\pi_{j}/\pi_{i}<g_{j}/g_{i}$ is more likely to occur than $\pi_{j}/\pi_{i}\geq g_{j}/g_{i}$, according to the effect. Since $\lambda_{ji}^{*}$ is between $g_{j}/g_{i}$ and $\pi_{j}/\pi_{i}$, the mitigation fills the gap between these two ratios without changing the likelihood.
Theorem 5
The Nash bargaining solution to Eq. (16), (17), or (18) is
$\left\{\begin{array}[]{rcl}\lambda_{ji}^{*}&\overset{\mathrm{def}}{=\joinrel=}&\sqrt{\frac{\pi_{j}g_{j}}{\pi_{i}g_{i}}},\hskip 14.22636pt\mathrm{and}\\ \lambda_{ij}^{*}&\overset{\mathrm{def}}{=\joinrel=}&\sqrt{\frac{\pi_{i}g_{i}}{\pi_{j}g_{j}}}\end{array}\right.$
for $\lambda_{ji}$ and $\lambda_{ij}$, respectively.
The solution of $\lambda_{ji}^{*}$ and $\lambda_{ij}^{*}$, in general, relies on the choice of status quo points. In Eq. (16), for example, both players could intensify or weaken their status quo positions. If the intensification or weakening is symmetrical with both players, we scale these two status quo points by a positive number $c\leq\sqrt{q/p}$. Then, $\lambda_{ji}$’s effective bargaining set extends from $[p,q]$ to $[cp,q/c]$ and $\lambda_{ij}$’s extends from $[1/q,1/p]$ to $[c/q,1/(cp)]$. However, the bargaining solution remains unchanged, according to Corollary 3. Letting $c\to 0$, the new effective bargaining sets become $(0,\infty)$ while the symmetry still bounds the solution of $\lambda_{ji}^{*}$ within $[p,q]$ and $\lambda_{ij}^{*}$ within $[1/q,1/p]$.
@@
For any positive $c\leq\sqrt{q/p}$, if we replace $p$ with $cp$ and $q$ with $q/c$ in Eq. (16), (17), or (18), the solution is still $\lambda_{ji}=\lambda_{ji}^{*}$ and $\lambda_{ij}=\lambda_{ij}^{*}$.
In summary, when country $i$ and $j$’s preferences are known, we recommend using the following $\lambda_{ji}$:
$\lambda_{ji}=\left\{\begin{array}[]{cl}\frac{\pi_{j}}{\pi_{i}},&\mathrm{if\ both\ prefer\ higher\ competitiveness;}\\ \frac{g_{j}}{g_{i}},&\mathrm{if\ both\ prefer\ better\ trade\ balance;}\\ \sqrt{\frac{\pi_{j}g_{j}}{\pi_{i}g_{i}}}&\mathrm{otherwise.}\end{array}\right.$ (19)
Otherwise, we suggest using $\lambda_{ji}=\sqrt{\pi_{j}g_{j}}/\sqrt{\pi_{i}g_{i}}$ as a rule of thumb. Country $j$ should avoid competing for a higher $\pi_{j}$ with a significantly larger-sized economy $i$, resulting in $\lambda_{ji}=\pi_{j}/\pi_{i}$. If it seeks a better trade balance instead, then the Nash solution expects $\lambda_{ji}=\sqrt{\pi_{j}g_{j}}/\sqrt{\pi_{i}g_{i}}$, which is likely larger than $\pi_{j}/\pi_{i}$ according to the Matthew effect. Consequently, by Eq. (10), $\pi_{j}$ rises as $j$’s exports to $i$ grow, no matter trade surplus or deficit. This was probably China’s strategy and practice from 1980 to 2010 when dealing with greater economic powers; Deng Xiaoping set its diplomatic policy to “keep a low profile and take no lead” (e.g., Yan, 2014), and trade policy is often used as a tool of foreign policy. Similarly, if country $j$ competes with $i$ for a better trade balance, then the solution $\lambda_{ji}=g_{j}/g_{i}>\pi_{j}/\pi_{i}$ and a decreasing $P_{ji}$ leads to a declining $\pi_{j}$ in Eq. (10). Thus, $j$ should also not compete with $i$. As a result, an even distribution of economic, political, and military capabilities between contending countries is likely to increase the chance of war, and peace is preserved best when there is an imbalance of national capabilities between them (e.g., Organski, (1968)’s power transition theory).
Finally, ${\pi_{j}}/{\pi_{i}}$ and ${g_{j}}/{g_{i}}$ in Eq. (19) are also Nash bargaining solutions for $\lambda_{ji}$ because
$\underset{\lambda_{ji}>0}{\mathrm{argmax}}\left\{\left(\lambda_{ji}-\frac{\pi_{j}}{\pi_{i}}\right)\left(\lambda_{ij}-\frac{\pi_{i}}{\pi_{j}}\right)\bigg{|}\lambda_{ij}\lambda_{ji}=1\right\}=\frac{\pi_{j}}{\pi_{i}}$ (20)
and
$\underset{\lambda_{ji}>0}{\mathrm{argmax}}\left\{\left(\lambda_{ji}-\frac{g_{j}}{g_{i}}\right)\left(\lambda_{ij}-\frac{g_{i}}{g_{j}}\right)\bigg{|}\lambda_{ij}\lambda_{ji}=1\right\}=\frac{g_{j}}{g_{i}}.$ (21)
In Eq. (20) and (21), both countries battle for better competitiveness and net trade balance, respectively. At the solution points, the objective functions are zero and no country gets extra welfare, already implied in Corollary 1 and Theorem 3.
4.3 Posterior Bargaining Power
In a real bargaining context, non-economic factors (such as military power, population, culture, tradition, legislation, and location) and alternative optimal criteria could considerably compromise the symmetry axiom and hence deviate the bargaining solutions substantially different from $\lambda_{ji}^{*}$ and $\lambda_{ij}^{*}$. Thus, the countries likely have unequal prior bargaining power in Theorem 5, and we introduce a posterior bargaining power that updates the prior one.
To extend the bilateral Nash solution in Theorem 5 to multilateral, we define
as country $i$’s posterior bargaining power or national capability in the trade system. It is normalized such that all bargaining power sums to one and $\beta_{i}$ can be compared over time. Also, it measures bargaining power in multinational negotiations. When allocating votes in forming or reforming a multinational alliance, for example, using the power could avoid any bilateral dialogs and discussions.
Clearly, $\lambda^{*}_{ij}=\beta_{i}/\beta_{j}$ for any $i,j\in\mathcal{N}$. The bargaining power $\{\beta_{i}\}_{i=1}^{n}$ construes a linear ordering for all countries and holds a consistent transitivity property for the Nash solution $\lambda_{ij}^{*}$. If $i$ has a negotiation advantage over $j$ and $j$ over $k$, i.e., $\beta_{i}>\beta_{j}$ and $\beta_{j}>\beta_{k}$, then $i$ necessarily has more bargaining power than $k$ since $\beta_{i}>\beta_{k}$. Specifically, $i$’s bargaining power over $k$ equals the ratio of that of $i$ over $j$ and that of $j$ over $k$, i.e., $\lambda_{ik}^{*}=\lambda_{ij}^{*}/\lambda_{jk}^{*}$ for any $j\in\mathcal{N}$.
Furthermore, there are two equally weighted and highly correlated determinants in $\beta_{i}$: the volume of production $g_{i}$ and the strength of competitiveness $\pi_{i}$. So, $\beta_{i}$ inherits some “controversial” properties from $\pi_{i}$. From the production aspect, $\beta_{i}$ declines when the companies in country $i$ rely on inputs from its partner countries. The more the companies depend on the partners, the less the country’s power to coerce concessions from the partners. From the consumption aspect, the bargaining power increases when the country consumes more imported goods and services by influencing the production of its counterparts. Between the determinants, $g_{i}$ is the country’s breadth of economy, and $\pi_{i}$ is predominantly the depth of the economy since the matrix $P$ in Eq. (3) already normalizes the production volumes. However, the positive correlation between them implies a risk of domino effect in which a decline of $g_{i}$ triggers a decrease of $\pi_{i}$, which incurs further reduction of $g_{i}$. Consequently, $\beta_{i}$ would shrink more than the initial drop of $g_{i}$.
5 Economic Globalization
Globalization refers to removing the barriers to the flow of financial products, goods, technology, information, and jobs across national borders and cultures. In the economic context, it describes the increasing interdependence of a country with the rest of the world, fostered through the trade of goods and services. A recent example is China’s entry into the WTO in 2001, while the United Kingdom’s Brexit in 2020 was a clear rejection of globalization. On the contrary, a protectionism measure could raise tariffs on all imported goods, set import quotas for all other countries, and enact stricter government regulations on all imports. In 2018, for example, the Trump administration imposed punishing tariffs on all steel and aluminum imports from all other countries (e.g., Washington Post, 2018).
Globalization creates winners and losers within a nation but it does not automatically compensate the losers. Therefore, a more inclusive globalization policy is needed, one that not only maximizes the winners’ profits but also considers the negative impacts on the losers. This section questions whether the country should retreat from global economic integration without directly attacking any specific country. Again, the objective is to maximize the country’s national competitiveness rather than the aggregate utility function of its trade firms. We consider actions that directly affect the rest of the world and neglect bilateral trade balances.
Country $i$’s globalization action triggers adjustments on the whole matrix $P$. Gaming on $P$, the country’s first move is on the diagonal element $P_{ii}$, say, with a small change $\Delta P_{ii}$ on $P_{ii}$. If it increases globalization, it shrinks its value of $P_{ii}$ and $\Delta P_{ii}<0$; otherwise, it increases $P_{ii}$ and $\Delta P_{ii}>0$ if it boosts protectionism. As the action has no specific target countries, we make the following assumption: whenever $P$ has a change in the $i$th column, the other columns change proportionally to ensure each row still has a unit sum. Thus, the change $\Delta P_{ii}$ on $P_{ii}$ provokes the adjustments $\Delta P$ on $P$ in three levels of fallout. First, for any $j\not=i$, $P_{ij}$’s proportional response to $\Delta P_{ii}$ is
$\Delta P_{ij}=-\frac{P_{ij}}{1-P_{ii}}\Delta P_{ii},$
and the $i$th row of $\Delta P$ is
$\Delta P_{ii}\left(-\frac{P_{i1}}{1-P_{ii}},\cdots,-\frac{P_{i,i-1}}{1-P_{ii}},1,-\frac{P_{i,i+1}}{1-P_{ii}},\cdots,-\frac{P_{in}}{1-P_{ii}}\right),$
which has a zero sum and the adjustments are proportional to the $i$th row of $P$. Secondly, the change $\Delta P_{ij}$ causes
$\Delta P_{ji}=\lambda_{ij}\Delta P_{ij}=-\frac{\lambda_{ij}P_{ij}}{1-P_{ii}}\Delta P_{ii}$
due to country $j$’s reprisal reaction. Therefore, the $i$th column of $\Delta P$ is
$\Delta P_{ii}\left(-\frac{\lambda_{i1}P_{i1}}{1-P_{ii}},\cdots,-\frac{\lambda_{i,i-1}P_{i,i-1}}{1-P_{ii}},1,-\frac{\lambda_{i,i+1}P_{i,i+1}}{1-P_{ii}},\cdots,-\frac{\lambda_{in}P_{in}}{1-P_{ii}}\right)^{\top}.$
Lastly, for any $t\not=i$, $P_{jt}$’s proportional response to $\Delta P_{ji}$ is
$\Delta P_{jt}=-\frac{P_{jt}}{1-P_{ji}}\Delta P_{ji}=\frac{\lambda_{ij}P_{ij}P_{jt}}{(1-P_{ii})(1-P_{ji})}\Delta P_{ii}.$
The sum of the $j$th row in $\Delta P$ is
$-\frac{\lambda_{ij}P_{ij}}{1-P_{ii}}\Delta P_{ii}+\sum\limits_{t\not=i}\frac{\lambda_{ij}P_{ij}P_{jt}}{(1-P_{ii})(1-P_{ji})}\Delta P_{ii}=\frac{\lambda_{ij}P_{ij}}{1-P_{ii}}\left[-1+\sum\limits_{t\not=i}\frac{P_{jt}}{1-P_{ji}}\right]\Delta P_{ii}=0.$
We define the matrix $M$ by
$\left[\begin{array}[]{ccccccc}\frac{\lambda_{i1}P_{i1}P_{11}}{1-P_{1i}}&\cdots&\frac{\lambda_{i1}P_{i1}P_{1,i-1}}{1-P_{1i}}&-\lambda_{i1}P_{i1}&\frac{\lambda_{i1}P_{i1}P_{1,i+1}}{1-P_{1i}}&\cdots&\frac{\lambda_{i1}P_{i1}P_{1n}}{1-P_{1i}}\\ \vdots&\ddots&\vdots&\vdots&\vdots&\ddots&\vdots\\ \frac{\lambda_{i,i-1}P_{i,i-1}P_{i-1,1}}{1-P_{i-1,i}}&\cdots&\frac{\lambda_{i,i-1}P_{i,i-1}P_{i-1,i-1}}{1-P_{i-1,i}}&-\lambda_{i,i-1}P_{i,i-1}&\frac{\lambda_{i,i-1}P_{i,i-1}P_{i-1,i+1}}{1-P_{i-1,i}}&\cdots&\frac{\lambda_{i,i-1}P_{i,i-1}P_{i-1,n}}{1-P_{i-1,i}}\\ -P_{i1}&\cdots&-P_{i,i-1}&1-P_{ii}&-P_{i,i+1}&\cdots&-P_{in}\\ \frac{\lambda_{i,i+1}P_{i,i+1}P_{i+1,1}}{1-P_{i+1,i}}&\cdots&\frac{\lambda_{i,i+1}P_{i,i+1}P_{i+1,i-1}}{1-P_{i+1,i}}&-\lambda_{i,i+1}P_{i,i+1}&\frac{\lambda_{i,i+1}P_{i,i+1}P_{i+1,i+1}}{1-P_{i+1,i}}&\cdots&\frac{\lambda_{i,i+1}P_{i,i+1}P_{i+1,n}}{1-P_{i+1,i}}\\ \vdots&\ddots&\vdots&\vdots&\vdots&\ddots&\vdots\\ \frac{\lambda_{in}P_{in}P_{n1}}{1-P_{ni}}&\cdots&\frac{\lambda_{in}P_{in}P_{n,i-1}}{1-P_{ni}}&-\lambda_{in}P_{in}&\frac{\lambda_{in}P_{in}P_{n,i+1}}{1-P_{ni}}&\cdots&\frac{\lambda_{in}P_{in}P_{nn}}{1-P_{ni}}\end{array}\right].$
Then $\Delta P=M\Delta P_{ii}/(1-P_{ii})$ and so, $\frac{\mathrm{d}P}{\mathrm{d}P_{ii}}=M/(1-P_{ii})$ as $\Delta P_{ii}\to 0$. In $M$, there are $22n-$ 2 negative numbers, all in the $i$th row or column; the others are positive. We also let $M_{i}$ be $M$ with its $i$th column dropped. Theorem 6 lists the derivative of $\pi$ with respect to $P_{ii}$.
Theorem 6
With the above setting and notations, we have
$\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=-\frac{1}{1-P_{ii}}\frac{\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}(\pi M_{i})^{\top}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}}$ (22)
and
$\frac{\mathrm{d}\pi_{-i}}{\mathrm{d}P_{ii}}=\frac{\left(I_{n-1}-Z_{i}\right)^{-1}}{1-P_{ii}}\left[\left(\pi M_{i}\right)^{\top}-\frac{\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}(\pi M_{i})^{\top}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}}\alpha_{i}\right].$ (23)
Thus, the decision rule on globalization could be like this: country $i$ would engage in further globalization or protectionism if $\frac{\mathrm{d}\log\pi_{i}}{\mathrm{d}\log P_{ii}}$ is significantly negative or positive, respectively. Otherwise, when $\frac{\mathrm{d}\log\pi_{i}}{\mathrm{d}\log P_{ii}}$ is insignificantly different from zero, country $i$ could seek targets of trade wars or trade partnerships as described in Sections 3 and 4. In between, it could form or participate in preferential trade agreements (PTAs) or regional trade agreements (RTA) with other economies. Within an ideal PTA or RTA, the ebb and flow of $\lambda_{j}/\lambda_{i}$ and $g_{j}/g_{i}$ should be well balanced to facilitate common agreements and to prevent major internal trade disputes. Also, location and cultural proximity often foster the formation of a PTA or RTA. Besides, Eq. (23) measures the side effects of $i$’s protectionism on the rest of the world. In general, some would benefit from it, whereas others would not.
Country $i$’s optimal level of globalization maximizes $\pi_{i}$ by adjusting $P_{ii}$ and, accordingly, rebalancing $P$; the first order condition is necessary at the optimal level. This can simply test if $\pi M_{i}$ is a zero vector or not, according to the first part of Corollary 4. Given $\lambda_{ji}$ for all $i$ and $j$ in $\mathcal{N}$, Eq. (22) and (23) for all $i\in\mathcal{N}$ define a dynamical system that describes the instant change of $\pi$ with respect to the diagonal of $P$. Given the system’s complexity, numerical algorithms may achieve a solution to $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}=\vec{0}_{n}^{\top}$ for all $i\in\mathcal{N}$ through successive tiny adjustments on $P$. In such an algorithm, for example, over-globalized countries increase $P_{ii}$ by 1% at each adjustment while under-globalized countries reduce $P_{ii}$ by 1% at the same time; and this succession continues until the matrix $P$ converges. At the potential convergence, the final $P$ and $\pi$ reach a new counterbalance equilibrium in Eq. (24). However, $P$ needs no more adjustments if $\Lambda$ satisfies certain desirable properties, such as balanced bilateral trade according to Theorem 3 and the second part of Corollary 4. Also, $P$ needs no adjustments if $\lambda_{ij}=\pi_{i}/\pi_{j}$ or all countries fight for competitiveness in every bilateral conflict. At the convergence, any tiny perturbation on $P_{ii}$ would have a nonsignificant effect on $\pi$, discouraging $i$ from making small changes. Thus, the diagonal of $P$ is relatively stable levels of globalization in these two special cases. Yet, even if a point satisfies all the first-order conditions, it is optimal for some countries only and not every country maximizes its $\pi_{i}$ at this point. Besides, the derivatives in Eq. (22) and (23) are based on the exogeneity assumption of $\Lambda$ and the proportional rebalancing in $M$. For these reasons, it might be of limited practical use to seek a far-reached solution to the first-order conditions, even if it exists. Nevertheless, Theorem 6 provides each country with a right direction to steer.
@@ 5 Economic Globalization
1. $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}=\vec{0}^{\top}_{n}$ if and only if $\pi M_{i}=\vec{0}^{\top}_{n-1}$; 2. If $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}=\vec{0}^{\top}_{n}$ for all $i\in\mathcal{N}$, then $\pi$ satisfies that
$\pi\left[\Lambda\bigotimes P\right]^{\top}=\pi$ (24)
where $\bigotimes$ is the element-wise matrix product; as two special cases, Eq. (24) is satisfied if $\lambda_{ij}=\pi_{i}/\pi_{j}$ for all $i$ and $j$ in $\mathcal{N}$, or $\lambda_{ij}=P_{ji}/P_{ij}$ for all $i$ and $j$ in $\mathcal{N}$.
Finally, Theorem 6 implies a mixed globalization and protectionism strategy. Let us divide the countries in $\mathcal{N}\setminus\{i\}$ into two groups based on their signs in the vector $\pi M_{i}$: country $j$ belongs to the group of $\mathcal{N}^{+}$ or $\mathcal{N}^{-}$ if its component in $\pi M_{i}$ is positive or negative, respectively. Because $(I_{n-1}-Z_{i})^{-1}$ is a non-negative matrix and $\alpha_{i}$ a non-negative vector, we rewrite Eq. (22) as $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=-(\pi M_{i})\omega$ for some non-negative vector $\omega$. Thus, when $j\in\mathcal{N}^{+}$ or $j\in\mathcal{N}^{-}$, it contributes negatively or positively, respectively, to $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}$; and the positive and negative contributions are partially offset. Given $\Delta P_{ii}$, we have the approximate change of $\pi_{i}$ as
$\Delta\pi_{i}\approx-\sum_{j\in\mathcal{N}^{+}}\left(\pi M_{i}\right)_{{}_{j}}\omega_{j}\Delta P_{ii}-\sum_{j\in\mathcal{N}^{-}}\left(\pi M_{i}\right)_{{}_{j}}\omega_{j}\Delta P_{ii}.$ (25)
In the mixed strategy, country $i$ engages in further protectionism with countries in $\mathcal{N}_{i}^{-}$ and further globalization with countries in $\mathcal{N}_{i}^{+}$, avoiding the partial offsetting. This strategy is likely better than a pure globalization or protectionism strategy because the right-hand side of Eq. (25) is less than or equal to
$-\sum\limits_{j\in\mathcal{N}^{+}}(\pi M_{i})_{{}_{j}}\omega_{j}(-|\Delta P_{ii}|)-\sum\limits_{j\in\mathcal{N}^{-}}(\pi M_{i})_{{}_{j}}\omega_{j}|\Delta P_{ii}|.$
Therefore, country $i$ should increase or decrease cooperation with countries in $\mathcal{N}^{+}$ or $\mathcal{N}^{-}$, respectively.

### v5

@@ 5 Economic Globalization
$\frac{\mathrm{d}\pi_{-i}}{\mathrm{d}P_{ii}}=\left(I_{n-1}-Z_{i}\right)^{-1}\left[\left(\pi M_{i}\right)^{\top}-\frac{\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}(\pi M_{i})^{\top}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}}\alpha_{i}\right].$ (18)
The optimal level of globalization for country $i$ maximizes $\pi_{i}$ by adjusting $P_{ii}$ and, accordingly, rebalancing $P$. The first-order condition $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=0$ is necessary at the optimal level because $P_{ii}$ cannot be $1$ or 00. However, $\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ii}}$ may be positive or negative at this optimal level, allowing country $j$ to capitalize on this by making a corresponding negative or positive change in $P_{ij}$, thereby affecting $P_{ii}$. To disincentive other countries from making simultaneous movements, we set $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}$ to be a zero vector. This can be tested by checking if $\pi M_{i}$ is a zero vector, according to Corollary 7.
@@
$\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}=\vec{0}^{\top}_{n}$ if and only if $\pi M_{i}=\vec{0}^{\top}_{n-1}$.
Given $\Lambda$, Eqs. (17) and (18) for all $i\in\mathcal{N}$ define a dynamical system describing the instantaneous change of $\pi$ with respect to the diagonal of $P$. There are $n$ independent variables in the system since the proportions among off-diagonal elements remain unchanged. Due to the system’s complexity, numerical algorithms may achieve a solution to $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}=\vec{0}_{n}^{\top}$ for all $i\in\mathcal{N}$ through successive tiny adjustments on the diagonal. For example, over-globalized countries increase $P_{ii}$ by 1% at each adjustment while under-globalized countries reduce $P_{ii}$ by 1% simultaneously; this succession continues until the matrix $P$ converges.
Alternatively, we study the properties of the solution to the first-order conditions. If the solution exists, it satisfies a new counterbalance equilibrium in Eq. (19), as stated in Corollary 8. Eq. (19) places $n-1$ identities to the solution, requiring only one more condition to uniquely identify it. Also, $P$ and $\pi$ already satisfy the new counterbalance if $\Lambda$ exhibits some desirable properties, such as $\lambda_{ij}=\pi_{i}/\pi_{j}$ or $P_{ji}/P_{ij}$, resulting in bilateral authority or trade balance, according to Eq. (6) or Theorem 3, respectively. At the stable solution, any tiny perturbation on the diagonal of $P$ would have a nonsignificant effect on $\pi$, discouraging countries from making small movement. Yet, this solution is optimal for some countries only, and not every country maximizes its competitiveness with this solution. Additionally, the derivatives in Eqs. (17) and (18) are based on the exogeneity assumption of $\Lambda$ and the proportional rebalancing in $M$. Therefore, seeking a far-reaching solution to the first-order conditions might be of limited practical use. Nevertheless, Theorem 6 provides each country with the right direction to steer.
@@
If $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}=\vec{0}^{\top}_{n}$ for all $i\in\mathcal{N}$, then $\pi$ satisfies that
$\pi\left[\Lambda\odot P\right]^{\top}=\pi.$ (19)
Eq. (19) is satisfied if $\lambda_{ij}=\pi_{i}/\pi_{j}$ or $\lambda_{ij}=P_{ji}/P_{ij}$ for all $i$ and $j$ in $\mathcal{N}$.
The decision rule on globalization could be as follows: country $i$ would engage in further globalization or protectionism if $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}$ is significantly negative or positive, respectively. Otherwise, when $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}$ is insignificantly different from zero, country $i$ could seek targets of trade wars or trade partnerships as described in Sections 3 and 4. In between, it could form or participate in preferential trade agreements (PTAs) or regional trade agreements (RTA) with other economies. Within an ideal PTA or RTA, the ebb and flow of $\lambda_{j}/\lambda_{i}$ and $g_{j}/g_{i}$ should be well balanced to facilitate common agreements and prevent major internal trade disputes. Location and cultural proximity often foster the formation of a PTA or RTA. Besides, Eq. (18) measures the side effects of $i$’s protectionism on the rest of the world, helping select countries for a PTA or RTA. In general, some would benefit from it, whereas others would not.
Finally, Theorem 6 implies a mixed globalization and protectionism strategy, helping $i$ to decide which PTA or RTA to join. Let us divide the countries in $\mathcal{N}\setminus\{i\}$ into two groups based on their signs in the vector $\pi M_{i}$: country $j$ belongs to the group of $\mathcal{N}_{i}^{+}$ or $\mathcal{N}_{i}^{-}$ if its component in $\pi M_{i}$ is positive or negative, respectively. Because $(I_{n-1}-Z_{i})^{-1}$ is a non-negative matrix and $\alpha_{i}$ a non-negative vector, we rewrite Eq. (17) as $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=-(\pi M_{i})\omega$ for some non-negative vector $\omega$. Thus, when $j\in\mathcal{N}_{i}^{+}$ or $j\in\mathcal{N}_{i}^{-}$, it contributes negatively or positively, respectively, to $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}$, and the positive and negative contributions are partially offset.
Given $\Delta P_{ii}$, we have the approximate change of $\pi_{i}$ as:
$\Delta\pi_{i}\approx-\sum_{j\in\mathcal{N}_{i}^{+}}\left(\pi M_{i}\right)_{{}_{j}}\omega_{j}\Delta P_{ii}-\sum_{j\in\mathcal{N}_{i}^{-}}\left(\pi M_{i}\right)_{{}_{j}}\omega_{j}\Delta P_{ii}.$ (20)
In the mixed strategy, country $i$ engages in further protectionism with countries in $\mathcal{N}_{i}^{-}$ and further globalization with countries in $\mathcal{N}_{i}^{+}$, avoiding the partial offsetting. This strategy is better than a pure globalization or protectionism strategy because the right-hand side of Eq. (20) is less than or equal to:
$-\sum\limits_{j\in\mathcal{N}_{i}^{+}}(\pi M_{i})_{{}_{j}}\omega_{j}(-|\Delta P_{ii}|)-\sum\limits_{j\in\mathcal{N}_{i}^{-}}(\pi M_{i})_{{}_{j}}\omega_{j}|\Delta P_{ii}|.$
Therefore, country $i$ should increase or decrease cooperation with countries in $\mathcal{N}_{i}^{+}$ or $\mathcal{N}_{i}^{-}$, respectively. If most country members of a PTA or RTA are in $\mathcal{N}_{i}^{+}$, then country $i$ should consider participating in the trade agreement. However, country $j\in\mathcal{N}_{i}^{+}$ does not necessarily imply that country $i\in\mathcal{N}_{j}^{+}$, possibly causing additional negotiations if $j$ is already in the PTA or RTA.